# From Sequential to Parallel: Reformulating Dynamic Programming as GPU Kernels for Large-Scale Stochastic Combinatorial Optimization

### v1

@@ 1 Introduction
We demonstrate the approach on two benchmark problems that capture core challenges in stochastic routing and inventory optimization: (i) a vectorized split operator for capacitated vehicle routing with stochastic demand (CVRPSD), and (ii) an inventory reinsertion DP under an order-up-to policy. Across benchmarks, our implementation scales nearly linearly with the number of scenarios and achieves substantial speedups over CPU baselines—from one–two orders of magnitude for CVRPSD to four–five orders for DSIRP. Additional experiments show that these gains translate directly into better decision quality: the ability to evaluate far more scenarios and far more first-stage candidates within the same time budget consistently yields stronger solutions. Our implementation is available at https://anonymous.4open.science/r/2-stage-IRP-GPU-0525 and https://anonymous.4open.science/r/2-stage-IRP-GPU-0516.
@@ 2 Generic Dynamic Programming Framework
The decision process follows a two-stage stochastic optimization framework. In the first stage, a delivery and routing plan is established for Day 1. Specifically, the model determines which customers to replenish and how much to deliver, subject to vehicle capacity and routing constraints. These decisions are made prior to the realization of demand and are identical across all demand scenarios. Once the demand over the entire planning horizon (Days 1 to $H$) is realized, second-stage decisions are made adaptively from Day 2 onward. These include scenario-dependent routing and replenishment actions that respond to the realized demand in each scenario (see the Figure 10 in A.3 for better understanding).
@@ A.1 Related Works
We consider a two-stage stochastic version of the Inventory Routing Problem (IRP) under an Order-Up-To (OU) inventory policy. We define the model on a complete directed graph $\mathcal{G}=(\mathcal{N},\mathcal{A})$, where $\mathcal{N}=\{0,n+1\}\cup\mathcal{N}^{\prime}$ includes the supplier node 0, the destination depot $n+1$, and the set of customers $\mathcal{N}^{\prime}$. The arc set $\mathcal{A}$ represents all possible directed connections between nodes. A set of vehicles $\mathcal{K}$ is available for deliveries, with $|\mathcal{K}|=K$. The planning horizon spans a finite set of days $\mathcal{T}=\{1,\ldots,H\}$ and $\mathcal{T^{\prime}}=\{2,\ldots,H\}$ for the second stage. To model uncertainty in demand, we consider a finite set of scenarios $\Omega$, where each scenario $\omega\in\Omega$ occurs with probability $p^{\omega}$. Each vehicle has a capacity limit $Q$, and each customer $i\in\mathcal{N}^{\prime}$ has an inventory capacity $U_{i}$. Holding costs $h_{i}$ are incurred per unit of inventory stored at node $i$, and a stock-out penalty is incurred at a rate of $\rho h_{i}$ per unit of unmet demand at customer $i$, where $\rho>1$. The customer demand at customer $i$ on day $t$ under scenario $\omega$ is denoted by $d_{i}^{t,\omega}$. The cost of traveling from node $i$ to node $j$ is denoted $c_{ij}$ for each arc $(i,j)\in\mathcal{A}$. Finally, $I_{i}^{0}$ represents the initial inventory level at node $i$ at the beginning of the planning horizon.
Figure 10: The decision process follows a two-stage stochastic optimization framework. In the first stage, a delivery and routing plan is established for Day 1. Specifically, the model determines which customers to replenish and how much to deliver, subject to vehicle capacity and routing constraints. These decisions are made prior to the realization of demand and are shared across all possible demand scenarios. Once the demand over the entire planning horizon (Days 1 to $H$) is realized, second-stage decisions are made adaptively from Day 2 onward. These include scenario-specific vehicle routing and replenishment actions that respond to the realized demand in each scenario. The objective is to minimize the total expected cost, which consists of two components: (1) First-stage costs, including the Day 1 routing cost and the supplier’s inventory holding cost; (2) Second-stage costs, which vary across scenarios and include: (i) Inventory holding costs and stock-out penalties at customers at the end of Day 1; (ii) Routing and delivery costs from Day 2 to $H$; (iii)Inventory holding and stock-out penalties from Day 2 to $H$.
To simplify replenishment decisions and reflect common logistics practices, we adopt an OU policy. Under this policy, each customer is either replenished up to its full capacity $U_{i}$, or not replenished at all on a given day. This is modeled using binary variables $z_{i}^{t}$ (or $z_{i}^{t,\omega}$ in the second stage), which equal 1 if customer $i$ is replenished on day $t$, and 0 otherwise.
Master Problem: First-Stage Optimization with Expected Future Cost: The first-stage of the 2SIRP involves determining the vehicle routing and delivery quantities on Day 1, before the actual demand realizations are observed. Specifically, the first-stage decisions include the quantity $q_{i}^{k,1}$ delivered to customer $i$ by vehicle $k$, as well as the routing and loading variables $z_{i}^{k,1},\ y_{ij}^{k,1},Q_{i}^{k,1}$ that describe the path and vehicle load throughout the tour. We use $x=\{q_{i}^{k,1},z_{i}^{k,1},\ y_{ij}^{k,1},Q_{i}^{k,1}\}$ to denote all the variables in this stage. Based on these decisions, we formulate the master problem to minimize the total cost on Day 1, including the supplier’s inventory holding cost and the routing cost, along with the expected second-stage cost-to-go $\tilde{Q}^{\omega}(x)$ across all demand scenarios $\omega\in\Omega$. The probability of scenario $\omega\in\Omega$ occurring is $p^{\omega}$ and $\sum_{\omega\in\Omega}p^{\omega}=1$.
$\min_{x}\quad$ $h_{0}I_{0}^{1}+\sum_{k\in\mathcal{K}}\sum_{(i,j)\in\mathcal{A}}c_{ij}y_{ij}^{k,1}+\sum_{\omega\in\Omega}p^{\omega}\tilde{Q}^{\omega}(x)$ (25)
s.t. $I_{0}^{1}=I_{0}^{0}+d_{0}^{1}-\sum_{k\in\mathcal{K}}\sum_{i\in\mathcal{N}^{\prime}}q_{i}^{k,1}$ (26)
$q_{i}^{k,1}=(U_{i}-I_{i}^{0})\cdot z_{i}^{k,1}$ $\forall i\in\mathcal{N}^{\prime},\ \forall k\in\mathcal{K}$ (27)
$\sum_{i\in\mathcal{N}^{\prime}}q_{i}^{k,1}\leq Q\cdot z_{0}^{k,1}$ $\forall k\in\mathcal{K}$ (28)
$\sum_{j\in\mathcal{N}^{\prime}\cup\{n+1\}}y_{ij}^{k,1}=z_{i}^{k,1},\quad\sum_{j\in\mathcal{N}^{\prime}\cup\{0\}}y_{ji}^{k,1}=z_{i}^{k,1}$ $\forall i\in\mathcal{N}^{\prime},\ \forall k\in\mathcal{K}$ (29)
$\sum_{j\in\mathcal{N}^{\prime}\cup\{n+1\}}y_{0j}^{k,1}=1,\quad\sum_{j\in\mathcal{N}^{\prime}\cup\{0\}}y_{j,n+1}^{k,1}=1$ $\forall k\in\mathcal{K}$ (30)
$Q_{0}^{k,1}=\sum_{i\in\mathcal{N}^{\prime}}q_{i}^{k,1}$ $\forall k\in\mathcal{K}$ (31)
$Q_{i}^{k,1}-q_{i}^{k,1}\geq Q_{j}^{k,1}-Q\cdot(1-y_{ij}^{k,1})$ $\forall(i,j)\in\mathcal{A},\ \forall k\in\mathcal{K}$ (32)
$q_{i}^{k,1},\ Q_{i}^{k,1},\ I_{0}^{1}\geq 0$ $\forall i\in\mathcal{N}^{\prime},\ \forall k\in\mathcal{K}$ (33)
$z_{i}^{k,1},\ y_{ij}^{k,1}\in\{0,1\}$ $\forall i\in\mathcal{N},\ (i,j)\in\mathcal{A},\ \forall k\in\mathcal{K}$ (34)
Constraint equation 26 tracks the inventory at the supplier after Day 1. Constraint equation 27 implements the OU policy: if a customer is visited by a vehicle, the vehicle must deliver exactly enough to fill the inventory up to its capacity $U_{i}$; otherwise, no delivery is made. Constraint equation 28 limits total delivery by vehicle capacity. Constraints equation 29–equation 30 ensure routing feasibility. Constraint equation 31 defines the starting load of each vehicle. Constraint equation 32 eliminates subtours. Finally, constraints equation 33–equation 34 define the variable domains.
Cost-to-Go Function and Second-Stage Problem Under Scenario $\omega$: Given the first-stage decision and realized demand $\bm{d}^{\omega}=\{d_{i}^{t,\omega}\}_{i\in\mathcal{N},t\in\mathcal{T^{\prime}}}$ under scenario $\omega$, the second-stage cost-to-go function $\tilde{Q}^{\omega}(x)$ can be calculated as: To ease our exposition, we will use the shorthand notations
$\bm{z}^{\omega}=\{z_{i}^{k,t,\omega}\}_{i\in\mathcal{N},k\in\mathcal{K},t\in\mathcal{T^{\prime}}}\quad$ $\bm{y}^{\omega}=\{y_{ij}^{k,t,\omega}\}_{(i,j)\in\mathcal{A},k\in\mathcal{K},t\in\mathcal{T^{\prime}}}$ (35)
$\bm{q}^{\omega}=\{q_{i,\omega}^{k,t}\}_{i\in\mathcal{N}^{\prime},k\in\mathcal{K},t\in\mathcal{T^{\prime}}}\quad$ $\bm{I}^{\omega}=\{I_{i,\omega}^{t}\}_{i\in\mathcal{N},t\in\mathcal{T^{\prime}}}\quad$ $\bm{B}^{\omega}=\{B_{i,\omega}^{t}\}_{i\in\mathcal{N}^{\prime},t\in\mathcal{T^{\prime}}}$ (36)
and use $\bm{\xi^{\omega}}=\{\bm{z^{\omega}},\bm{y^{\omega}},\bm{q}^{\omega},\bm{I}^{\omega},\bm{B}^{\omega}\}$, $\omega\in\Omega$ to denote all the variables in the second stage. The objective for the second stage can be represented by
$\tilde{Q}^{\omega}(x)=\sum_{i\in\mathcal{N^{\prime}}}\left(h_{i}I_{i}^{1,\omega}+\rho h_{i}B_{i}^{1,\omega}\right)+\min_{\bm{\xi^{\omega}}}\sum_{t=2}^{H}\bigg[h_{0}I_{0}^{t,\omega}+\sum_{i\in\mathcal{N^{\prime}}}\left(h_{i}I_{i}^{t,\omega}+\rho h_{i}B_{i}^{t,\omega}\right)+\sum_{k\in\mathcal{K}}\sum_{(i,j)\in\mathcal{A}}c_{ij}y_{ij}^{k,t,\omega}\bigg]$
s.t. $I_{i}^{1,\omega}=I_{i}^{0}+\sum_{k\in\mathcal{K}}q_{i}^{k,1}-d_{i}^{1,\omega}+B_{i}^{1,\omega}$ $\forall i\in\mathcal{N^{\prime}}$ (37)
$I_{i}^{t,\omega}=I_{i}^{t-1,\omega}+\sum_{k\in\mathcal{K}}q_{i}^{k,t,\omega}-d_{i}^{t,\omega}+B_{i}^{t,\omega}$ $\forall i\in\mathcal{N^{\prime}},\ t\in\mathcal{T^{\prime}},\omega\in\Omega$ (38)
$\sum_{k\in\mathcal{K}}q_{i}^{k,t,\omega}=(U_{i}-I_{i}^{t-1,\omega})\cdot z_{i}^{t,\omega}$ $\forall i\in\mathcal{N^{\prime}},\ t\in\mathcal{T^{\prime}},\omega\in\Omega$ (39)
$\sum_{i\in\mathcal{N^{\prime}}}q_{i}^{k,t,\omega}\leq Q\cdot z_{0}^{k,t,\omega}$ $\forall k\in\mathcal{K},\ t\in\mathcal{T^{\prime}},\omega\in\Omega$ (40)
$\sum_{j\in\mathcal{N^{\prime}}\cup{n+1}}y_{ij}^{k,t,\omega}=z_{i}^{t,\omega},\quad\sum_{j\in\mathcal{N^{\prime}}\cup{0}}y_{ji}^{k,t,\omega}=z_{i}^{t,\omega}$ $\forall i\in\mathcal{N^{\prime}},\ k\in\mathcal{K},\ t\in\mathcal{T^{\prime}},\omega\in\Omega$ (41)
$\sum_{j\in\mathcal{N^{\prime}}\cup{n+1}}y_{0j}^{k,t,\omega}=1,\quad\sum_{j\in\mathcal{N^{\prime}}\cup{0}}y_{j,n+1}^{k,t,\omega}=1$ $\forall k\in\mathcal{K},\ t\in\mathcal{T^{\prime}},\omega\in\Omega$ (42)
$Q_{0}^{k,t,\omega}=\sum_{i\in\mathcal{N^{\prime}}}q_{i}^{k,t,\omega}$ $\forall k\in\mathcal{K},\ t\in\mathcal{T^{\prime}},\omega\in\Omega$ (43)
$Q_{i}^{k,t,\omega}-q_{i}^{k,t,\omega}\geq Q_{j}^{k,t,\omega}-Q(1-y_{ij}^{k,t,\omega})$ $\forall(i,j)\in\mathcal{A},\ k\in\mathcal{K},\ t\in\mathcal{T^{\prime}},\omega\in\Omega$ (44)
$I_{i}^{t,\omega},\ B_{i}^{t,\omega},\ q_{i}^{k,t,\omega},\ Q_{i}^{k,t,\omega}\geq 0$ $\forall i\in\mathcal{N^{\prime}},\ k\in\mathcal{K},\ t\in\mathcal{T^{\prime}},\omega\in\Omega$ (45)
$z_{i}^{t,\omega},\ y_{ij}^{k,t,\omega}\in\{0,1\}$ $\forall i\in\mathcal{N},\ (i,j)\in\mathcal{A},\ k\in\mathcal{K},\ t\in\mathcal{T^{\prime}},\omega\in\Omega$ (46)
The second-stage constraints under the OU policy govern the evolution of inventory, replenishment quantities, and routing decisions from Day 2 to Day $H$ under each demand scenario $\omega\in\Omega$. Constraint equation 37 sets the initial inventory level for each customer $i\in\mathcal{N^{\prime}}$ at the end of Day 1, based on the first-stage deliveries, realized demand $d_{i}^{1,\omega}$, and any unmet demand carried as backorders $B_{i}^{1,\omega}$. Constraint equation 38 recursively updates the inventory and backorders for subsequent days $t\geq 2$ using the delivery quantities and realized daily demands. Constraint equation 39 enforces the OU policy: a customer either receives a shipment that raises its inventory up to the full capacity $U_{i}$ (i.e., $q_{i}^{k,t,\omega}=U_{i}-I_{i}^{t-1,\omega}$), or receives nothing. The binary variable $z_{i}^{t,\omega}$ indicates whether a replenishment occurs at node $i$ on day $t$. Constraint equation 40 ensures that the total quantity delivered by any vehicle $k$ on day $t$ does not exceed its capacity $Q$. Constraints equation 41 maintain routing feasibility: if a customer is visited ($z_{i}^{t,\omega}=1$), then exactly one arc must enter and one must leave it for each vehicle $k$. Constraint equation 42 ensures that each vehicle departs from the origin depot and ends at the destination depot once per day. Constraint equation 43 defines the load carried by each vehicle upon departure, equal to the sum of deliveries assigned to it. Constraint equation 44 eliminates subtours by imposing consistency in vehicle loads across successive arcs in a tour. Finally, constraints equation 45–equation 46 impose the appropriate domain restrictions, ensuring non-negativity of inventory, deliveries, and vehicle loads, and binary decisions for vehicle routing and customer visits.
Illustrative Example of the problem.
Figure 11: Illustration of the dynamic programming (DP) routine under the Order-Up-To (OU) policy. Blue bars represent possible inventory states, network diagrams show routing decisions, and arrows indicate state transitions with their associated costs. Infeasible transitions are marked as $\infty$. The DP systematically propagates costs forward, capturing the before-and-after evolution of replenishment paths; this routine can be embedded within any heuristic or metaheuristic framework.
The proposed framework is modular: while the outer layer can be any heuristic or metaheuristic search procedure that explores alternative first-stage decisions (e.g., different delivery plans or routing structures), the inner DP routine ensures that inventory evolution and replenishment actions are evaluated consistently across time and scenarios. Figure 11 illustrates how DP propagates costs forward by updating inventory states and associated routing paths under the OU policy. The blue bars denote customer inventory states, the network diagrams represent feasible routing actions, and the arrows indicate transitions with their corresponding costs. Infeasible transitions are marked as $\infty$. By capturing the before-and-after evolution of replenishment paths, the DP routine provides a structured way to evaluate candidate solutions, regardless of the outer search framework in which it is embedded.

### v2

@@ 1 Introduction
We demonstrate the approach on two benchmark problems that capture core challenges in stochastic routing and inventory optimization: (i) a vectorized split operator for capacitated vehicle routing with stochastic demand (CVRPSD), and (ii) an inventory reinsertion DP under an order-up-to policy. Across benchmarks, our implementation scales nearly linearly with the number of scenarios and achieves substantial speedups over CPU baselines—from one–two orders of magnitude for CVRPSD to four–five orders for DSIRP. Additional experiments show that these gains translate directly into better decision quality: the ability to evaluate far more scenarios and far more first-stage candidates within the same time budget consistently yields stronger solutions. Our implementation is available at https://github.com/Jingyi-poly/2-stage-IRP-GPU/tree/CVRPSD-split-GPU and https://github.com/Jingyi-poly/2-stage-IRP-GPU/tree/OUIRP-GPU.
@@ 2 Generic Dynamic Programming Framework
The decision process follows a two-stage stochastic optimization framework. In the first stage, a delivery and routing plan is established for Day 1. Specifically, the model determines which customers to replenish and how much to deliver, subject to vehicle capacity and routing constraints. These decisions are made prior to the realization of demand and are identical across all demand scenarios. Once the demand over the entire planning horizon (Days 1 to $H$) is realized, second-stage decisions are made adaptively from Day 2 onward. These include scenario-dependent routing and replenishment actions that respond to the realized demand in each scenario (see the Figure LABEL:deter in A.3 for better understanding).
@@ 4 Conclusion
Acknowledgments
This work was supported by the Longgang District Special Funds for Science and Technology Innovation under Grant LGKCSDPT2025002.
@@ A.1 Related Works
We consider a two-stage stochastic version of the Inventory Routing Problem (IRP) under an Order-Up-To (OU) inventory policy. We define the model on a complete directed graph $\mathcal{G}=(\mathcal{N},\mathcal{A})$, where $\mathcal{N}=\{0,n+1\}\cup\mathcal{N}^{\prime}$ includes the supplier node 0, the destination depot $n+1$, and the set of customers $\mathcal{N}^{\prime}$. The arc set $\mathcal{A}$ represents all possible directed connections between nodes. A set of vehicles $\mathcal{K}$ is available for deliveries, with $|\mathcal{K}|=K$. The planning horizon spans a finite set of days $\mathcal{T}=\{1,\ldots,H\}$ and $\mathcal{T^{\prime}}=\{2,\ldots,H\}$ for the second stage. To model uncertainty in demand, we consider a finite set of scenarios $\Omega$, where each scenario $\omega\in\Omega$ occurs with probability $p^{\omega}$. Each vehicle has a capacity limit $Q$, and each customer $i\in\mathcal{N}^{\prime}$ has an inventory capacity $U_{i}$. Holding costs $h_{i}$ are incurred per unit of inventory stored at node $i$, and a stock-out penalty is incurred at a rate of $\rho h_{i}$ per unit of unmet demand at customer $i$, where $\rho>1$. The customer demand at customer $i$ on day $t$ under scenario $\omega$ is denoted by $d_{i}^{t,\omega}$. The cost of traveling from node $i$ to node $j$ is denoted $c_{ij}$ for each arc $(i,j)\in\mathcal{A}$. Finally, $I_{i}^{0}$ represents the initial inventory level at node $i$ at the beginning of the planning horizon. To simplify replenishment decisions and reflect common logistics practices, we adopt an OU policy. Under this policy, each customer is either replenished up to its full capacity $U_{i}$, or not replenished at all on a given day. This is modeled using binary variables $z_{i}^{t}$ (or $z_{i}^{t,\omega}$ in the second stage), which equal 1 if customer $i$ is replenished on day $t$, and 0 otherwise.
The first-stage of the 2SIRP involves determining the vehicle routing and delivery quantities on Day 1, before the actual demand realizations are observed. Given the first-stage decision and realized demand $\bm{d}^{\omega}=\{d_{i}^{t,\omega}\}_{i\in\mathcal{N},t\in\mathcal{T^{\prime}}}$ under scenario $\omega$, the second-stage cost function $\tilde{Q}^{\omega}(x)$ can be calculated as $h_{0}I_{0}^{1}+\sum_{k\in\mathcal{K}}\sum_{(i,j)\in\mathcal{A}}c_{ij}y_{ij}^{k,1}+\sum_{\omega\in\Omega}p^{\omega}\tilde{Q}^{\omega}(x)$. The second-stage constraints under the OU policy govern the evolution of inventory, replenishment quantities, and routing decisions from Day 2 to Day $H$ under each demand scenario $\omega\in\Omega$. Constraints enforce the OU policy: a customer either receives a shipment that raises its inventory up to the full capacity $U_{i}$ (i.e., $q_{i}^{k,t,\omega}=U_{i}-I_{i}^{t-1,\omega}$), or receives nothing. Constraints also need to ensure cother classical IRP constraints, such as the total quantity delivered by any vehicle $k$ on day $t$ does not exceed its capacity $Q$.